Hydrogen sulfide reacts with oxygen to form sulfur dioxide and water according to the following unbalanced reaction: H2S (g) + O2 (g) —> SO2 (g) + H2O(g)
If we start with 36.6 g of Hudrogen Sulfide and 15.4 of oxygen, how many grams of water do we make?
1
Expert's answer
2020-11-25T02:20:49-0500

The mole ratio for the reaction="1:1"

Moles of water produced from"H_2S=" "32gH_2S\times" "1mol H_2S\over 34.1gH_2S" "\times" "1molH_2O\over 1mol H_2S" "\times" "18gH_2O\over 1mol H_2O" "=16.94gH_2O"

Moles of water produced from "O_2="

"15.4gH_2O\times" "1mol O_2\over 16.0gH_2O" "\times" "1mol H_2O\over 1mol O_2" "\times" "18gH_2O\over 1molH_2O" "=17.33gH_2O"

Total grams produced in the reaction"=16.94g+17.33g=34.27gH_2O"